A 1500.OmL container holds 7.80g of a gas at 2.35 atm and 25.00 C. How many moles of the gas are in the container ?


1
Expert's answer
2021-03-04T06:38:01-0500

According to the ideal gas law:

pV = nRT

where p - pressure (2.35 atm), V - volume (1500.0 ml = 1.50 L), n - number of moles, R - gas constant (0.082 L atm mol-1 K-1), T - temperature (25.00 °C = 298.15 K).

From here:

n = pV / RT = (2.35 atm × 1.50 L) / (0.082 atm mol-1 K-1 × 298.15 K) = 0.14 mol


Answer: 0.14 mol